# Dynamics of Clusters of Anyons in Fractional Quantum Hall Fluids

### v1

@@ II.3 Interaction Within Three Laughlin 1/3 Anyon Cluster
The interaction of a three-anyon cluster induced by realistic e-e interactions, including Coulomb and Yukawa interaction with different $\lambda$, is also studied. With LLL Coulomb interaction, the three-anyon cluster’s lowest energy state is $\Delta M=12$ (Figure 2 (a)), indicating a quite loosely bound structure. However, the configuration transitions to a more compact cluster with $\Delta M=6$ as $\lambda$ in the Yukawa potential increases (Figure 2 (b)), corresponding to shorter-range electron-electron interactions. This transition is shown by depicting the energy difference between $\Delta M=6$ ($E_{\Delta M=6}$) and its competing states $\Delta M=8$ or $\Delta M=12$ as a function of $\lambda$ in Figure 2 (c). $\Delta M=6$ gradually becomes the lowest energy state as $\lambda$ increases. This aligns with our earlier finding in two-anyon cases that sufficiently short-range interactions dominated by $\hat{V}_{1}$ can replicate the effects of the model Hamiltonian Eq.(8) on the effective anyon interaction. It can also be seen from the whole spectrum of the system that the quasihole excitations (red crosses) gradually separate from other states (blue pluses) when increasing the $\lambda$. We estimate the size of the anyon cluster using the first and second moments of their density distribution Johri et al. (2014). The estimated spatial radius of the lowest-energy anyon cluster is $40\sim 60\text{nm}$. See Supplementary Materials for more details.
@@ III Interactions within a cluster of Moore-Read Anyons
The Moore-Read (MR) FQH fluid has attracted much greater attention because of its non-Abelian nature and potential applications in quantum computations. Unlike the case of the Laughlin state, two quasiholes each with a $e/4$ charge will be created by one single flux insertion in the MR ground state because of the flux fractionalization in non-Abelian phases. When fused together, the two-quasihole excitation with $e/2$ charge collectively carries topological quantum numbers “$1$” or “$\psi$”, exhibiting a statistical phase 00 or $\pi$ respectively Baraban et al. (2009). The topological numbers of these two kinds of excitations come from the fact that they follow the Ising model in conformal field theory (CFT) with energetically degenerate fusion channels “$1$” and “$\psi$”, corresponding to the U(1) boson and Majorana fermion chiral edge modes Read and Moore (1991); Nayak and Wilczek (1996); Nayak et al. (2008); Möller et al. (2011). With the same reason, a single $e/4$ quasihole excitation corresponds to the index “$\sigma$” in Ising CFT model. In a general MR phase, the “$1$” and “$\psi$” types of quasihole excitations coexist, which is the reason for non-Abelian statistics. The degeneracy of a MR phase containing $N_{qh}$ quasiholes is $2^{N_{qh}/2-1}$.
@@ III.1 Dynamics of Two MR Anyons
The root configuration for the MR ground state is $110011...0011$, satisfying the $(4,2)$ admissible rule and the condition $N_{o}=2N_{e}-2$. We begin with the case of two quasiholes (one extra flux quanta) where the topological number of the system can only be either 1 or $\psi$, depending on whether the number of electrons ($N_{e}$) in the finite system is even or odd. These two types of two-quasihole excitations are illustrated by the following root configurations:
[…]
@@
First, the effective interaction between two MR anyons under the model e-e interaction is studied. Haldane pseudopotentials $\hat{V}_{1}$ and $\hat{V}_{3}$ are diagonalized within the null space of the MR model Hamiltonian, defined by a special three-body interactionGreiter et al. (1992); Simon et al. (2007), with the presence of two quasiholes. The effective anyon interactions defined by Eq.(9) with $N_{qh}=2$ are shown in Figure 3 (a) and (b), plotted against their relative angular momentum $\Delta M$. It depends on the parity of the system, as a manifestation of the odd-even effect in the MR phase Baraban et al. (2009). Since the resulting dynamics have distinct behaviors under different PPs by which the e-e interaction can always be decomposed, it suggests that tailored electron-electron interactions could exploit this distinction to manipulate the two types of MR anyons. For instance, the e-e interaction dominated by $\hat{V}_{1}$ induces the effective interactions for $\psi$-type anyons to bind together while favoring a comparatively large structure of $1$-types. Conversely, to introduce an effectively repulsive interaction between the $\psi$-type anyons for non-abelian braiding (therefore the possibility of topological quantum computation), e-e interaction dominated by $\hat{V}_{3}$ is preferred.
@@
The resultant effective interactions between two localized $e/4$ MR quasiholes are plotted as a function of their chord distance. The odd-even effect remains prominent with nearly opposite dynamical characteristics observed between systems with the two different parities. Under Coulomb interaction shown in Figure 3 (c), the effective interactions are repulsive at all distances for systems with $N_{e}$ of both parities. With shorter-range Yukawa interactions-for example $\lambda=0.5/l_{B}$ in Figure 3 (d)-the effective interaction for $\psi$-types is attractive at small separations. But it remains repulsive for $1$-types with the same distances then developing to be attractive at a comparatively large separation. This behaviour indicates that short-range interactions energetically favor tightly bound pairs of $e/4$ MR anyons in odd $N_{e}$ systems, while promoting spatially extended, loose configurations of them in even $N_{e}$ systems. It means that the $\psi$ fusion channel is energetically more favorable under the short-range interaction than the $1$ channel when considering the fusion of two “$\sigma$” anyons with $e/4$ charges. These results further support the feasibility of engineering bare electron-electron interactions to selectively separate and control the two types of fusion channels of the $e/4$ MR anyon and the consequent $e/2$ excitations.
@@ III.2 Interaction Within Four MR Anyon Clusters
We now look at a cluster of four MR quasiholes—each carrying a $e/4$ charge—which will be useful in Sec. IV. A four-MR-anyon cluster emerges upon the removal of a single electron from the ground state. The resulting system exhibits a degeneracy of $22^{4/2}-1=$ 2, where $1$- and $\psi$-types could coexist, adding significant complexity to the analysis. Specifically, the two pairs of quasiholes can only be two 1-types or two $\psi$-types in even $N_{e}$ systems, whereas the configuration necessarily consists of one 1-type and one $\psi$-type in odd $N_{e}$ systems Trung et al. (2025). The energy spectrum of this system under model electron-electron interactions $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ in the MR null space (Figure S3) reveals two key features: a persistent odd-even effect in systems with four MR quasiholes and a preference of loose anyon configurations characterized by large relative angular momentum $\Delta M$ within the four-anyon cluster.
To extend our investigation to more realistic scenarios, given that MR states are proposed to be stabilized in the 1LL, we studied the e-e Coulomb and Yukawa potentials in the 1LL shown in Figure 4. In contrast to Laughlin’s case, the ground states of MR systems with four anyons consistently exhibit a loose structure and remain largely insensitive to variations in electron-electron interactions. For the system with $N_{e}=13$, the lowest energy state is the one with relative angular momentum $\Delta M=12$, the maximum allowed $\Delta M$ value in this finite system. In the $N_{e}=14$ case, the four-anyon cluster attains the largest possible value of $\Delta M=14$ under the 1LL electron-electron Coulomb interaction, and changes to $\Delta M=12$ for both Yukawa interactions we computed with $\lambda=0.5$ and $1$. The results are not conclusive due to the limited system size we can reach. Another observation is that the quasihole excitation spectrum maintains a comparatively robust, model-independent structure, unchanged not only across the parameter regimes studied among the realistic interactions, but also across the realistic interaction and the model Hamiltonian inside the MR CHS, whose spectrum can be further found in Supplementary Materials. Using the first and second moments of the density distribution, the estimated real-space radius of the four-quasihole cluster is $100\sim 140\text{nm}$ for the lowest energy state as $\Delta M=12$ in $N_{e}=15$ (odd) system. And for the $\Delta M=14$ anyon cluster in $N_{e}=16$ (even) system, it is $75\sim 110\text{nm}$.
@@ IV Bulk tunneling experiments
The effective interactions within anyon clusters have direct experimental relevance in recently developed STM measurements, whose detailed implementations have been discussed in several recent works Papić et al. (2018); Hu et al. (2023); Coissard et al. (2022). In these experiments, single-electron tunneling occurs when the electronic potential $V_{t}$ applied to the tip satisfies
@@
where $E_{0}$ is the work function of the 2DEG, $E_{z}$ accounts for the finite-thickness effect Peterson et al. (2008), and $E_{xy}$ represents the energy scale of the strongly correlated 2D system studied here. After the bulk tunneling, a few-anyon cluster with total positive charge of $+e$ will be left behind. The number of anyons in this cluster is determined by the specific FQH phase (e.g. $3$ for Laughlin 1/3 and $4$ for MR). Due to the non-trivial inter-anyonic interactions as we have shown in the previous sections, the anyon cluster could be in different forms of “bound states” depending on the effective anyon interaction, the size of the STM tip, and the bias voltage applied to the tip. In particular, the effective interactions of such few-anyon clusters significantly influence the experimental outcomes, making such measurements potentially very useful for understanding the dynamics of anyon clusters.
@@
Figure 5: Possible total low-energy LDOS for Laughlin 1/3 state with $N_{e}=10$ with different resolutions (FWHM in the graphs). The entire background energies are subtracted in all the figures since only the energy differences matter here. (a) Coulomb interaction, (b) Yukawa interaction when $\lambda=1$, both in the LLL.
@@
The local density of states (LDOS) is a key observable in STM measurements. High-resolution tunneling can reveal a clear gap of approximately $0.1meV$, which corresponds to $\Delta E=0.002\sim 0.006\ e^{2}/\epsilon l_{B}$ whose exact value depends on the specific dielectric constant $\epsilon$ and the magnetic field $B$ under detection. Using several values within this range to be the full width at half-maximum (FWHM), we simulated the possible LDOS for the 3-anyon Laughlin 1/3 states and 4-anyon MR states. It is found that the variations in effective anyon interactions—induced by different bare e-e interactions—lead to distinct LDOS patterns, including differences in the numbers and positions of the peaks. Previous works mainly focus on the tunneling involving the most tightly bound anyon clusters Pu et al. (2024); Gattu et al. (2024). Our analysis show that the tunneling spectrum is very rich, and the most tightly bound anyon clusters both in the Laughlin and Moore-Read phases only occur at relatively higher energy (namely with higher bias voltage at the tip) that may merge with gapped excitations in these topological phases.
@@
We only present a preliminary study here, assuming the bias voltage of the STM tip does not significantly alter the density of the states of the anyon clusters. This may not be the case given the one-body potential highly depends on the geometry of the tip, the bias voltage, as well as the distance between the tip and the Hall manifold Trung et al. (2025). A more detailed study with explicit experimental parameters will be presented in the future.
@@ IV.1 LDOS Measurement for Laughlin 1/3 Phase
The low-energy spectrum for a system containing three Laughlin 1/3 anyons with realistic electron-electron interactions has already been shown in Figure 2 (a). The most tightly bound anyon cluster state ($\Delta M=0$ state) lies at significantly higher energy than the other low-lying excitations. The simulated LDOS for FWHM $0.003$ and $0.005$ under e-e Coulomb interaction and $\lambda=0.5$ Yukawa interaction both in LLL are shown in Figure 5. For relatively long-range e-e interactions such as LLL Coulomb, this state is buried among higher-energy excitations, making it undetectable as an isolated peak in electron tunneling. As the screening parameter $\lambda$ increases, all quasihole excitations shift downward and gradually separate from the high-energy continuum. The $\Delta M=0$ state eventually becomes distinguishable from the gapped excitations but remains the highest in energy among all quasihole states. The LDOS results for different interactions are clearly distinguishable, as seen by comparing (a) and (b) in Figure 5. This indicates that LDOS can serve as a fingerprint for characterising the effective interactions among anyons.
Figure 6: LDOS of the MR system containing 4 anyons for: 1LL Coulomb interaction with (a) $N_{e}=13$ and (b) $N_{e}=14$; 1LL Yukawa interaction when $\lambda=0.5$ with (c) $N_{e}=13$ and (d) $N_{e}=14$.

### v2

@@ II.3 Interaction Within Three Laughlin 1/3 Anyon Cluster
We studied the three-anyon spectrum for both Coulomb and Yukawa potentials with varying screening length $\lambda$. In all panels of Figure 2, red crosses mark states in the quasihole CHS, identified via a large-t projection using the model Hamiltonian $\hat{H}=t\hat{V}_{1}+\hat{V}_{\text{int}}$, where there exists a finite energy gap between the quasihole manifold and the other higher-energy states. They are tracked as $t\rightarrow 0$ toward the physical interaction, during which the structure of the QH manifold – particularly its state counting – remains unchanged because of its topological nature [69]. Blue plus symbols denote all the other states outside the quasihole manifold, which are not the subject of this paper and won’t be further identified.
With LLL Coulomb interaction, the three-anyon cluster’s lowest energy state is $\Delta M=12$, indicating a loosely bound structure. However, the configuration transitions to a more compact cluster with $\Delta M=6$ as $\lambda$ in the Yukawa potential increases, corresponding to shorter-range electron-electron interactions. This transition is shown by depicting the energy difference between $\Delta M=6$ ($E_{\Delta M=6}$) and its competing states $\Delta M=8$ or $\Delta M=12$ as a function of $\lambda$ in Figure 2 (c). $\Delta M=6$ gradually becomes the lowest energy state as $\lambda$ increases. This aligns with our earlier finding in two-anyon cases that sufficiently short-range interactions dominated by $\hat{V}_{1}$ can replicate the effects of the model Hamiltonian Eq.(9) on the effective anyon interaction. It can also be seen from the whole spectrum of the system that the quasihole excitations (red crosses) gradually separate from other states (blue pluses) when increasing the $\lambda$, since during this process, $\hat{V}_{1}$ gradually becomes more dominant. We estimate the size of the anyon cluster using the first and second moments of their density distribution [21]. The estimated spatial radius of the lowest-energy anyon cluster is $40\sim 60\text{nm}$. Density for the system under Coulomb e-e interaction indicates long-range correlations and oscillatory structures, while for Yukawa interaction, it has a prominent central peak and mild, smooth oscillations extending toward the edges, reflecting strong short-range binding  [70].
@@ III Interactions within a cluster of Moore-Read Anyons
The Moore-Read (MR) FQH fluid has attracted much greater attention because of its non-Abelian nature and potential applications in quantum computations. Unlike the case of the Laughlin state, two quasiholes, each with a $e/4$ charge, will be created by one single flux insertion in the MR ground state because of the flux fractionalization in non-Abelian phases. When fused, the two-quasihole excitation with $e/2$ charge collectively carries topological quantum numbers “11” or $\psi$”, exhibiting a statistical phase 0 or $\pi$ respectively [17]. The topological numbers of these two kinds of excitations come from the fact that they follow the Ising model in conformal field theory (CFT) with energetically degenerate fusion channels “11” and $\psi$”, corresponding to the U(1) boson and Majorana fermion chiral edge modes [4, 71, 15, 72]. For the same reason, a single $e/4$ quasihole excitation corresponds to the index $\sigma$” in the Ising CFT model. In a general MR phase, the “11” and $\psi$” types of quasihole excitations coexist, which is the reason for non-Abelian statistics. The degeneracy of a MR phase containing $N_{qh}$ quasiholes is $2^{N_{qh}/2-1}$.
@@ III.1 Dynamics of Two MR Anyons
The root configuration for the MR ground state is 1100110011110011...0011, satisfying the $(4,2)$ admissible rule and the condition $N_{o}=2N_{e}-2$. We begin with the case of two quasiholes (one extra flux quanta) where the topological number of the system can only be either 1 or $\psi$, depending on whether the number of electrons ($N_{e}$) in the finite system is even or odd. These two types of two-quasihole excitations are illustrated by the following root configurations:
 1 $\hbox{\hskip 6.66664pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-6.66664pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ\circ$}}\hskip-6.66664pt\hskip 6.66664pt}01100110011...110011$ 11-type, even $N_{e}$ (12)
 1 $\hbox{\hskip 6.66664pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-6.66664pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ\circ$}}\hskip-6.66664pt\hskip 6.66664pt}00110011...11001$1 $\psi$-type, odd $N_{e}$ (13)
@@
To compute the self-energy of a single $e/4$ quasihole in the MR phase, we fix the two quasiholes at the north and south poles of the sphere by Jack polynomials as 1010101011010...10101 and extract the energy from the ground-state energy $E_{g}$ with the same electron number. In the large-system limit, where their mutual interaction vanishes, the residual energy should approach twice the quasihole self-energy. More detailed explanations and discussion are provided in the Supplementary Material.
@@
First, the effective interaction between two MR anyons under the model e-e interaction is studied. Haldane pseudopotentials $\hat{V}_{1}$ and $\hat{V}_{3}$ are diagonalized within the null space of the MR model Hamiltonian, defined by a special three-body interaction[73, 24] and obtained by a similar strategy as Laughlin’s case Eq.(9) numerically, with the presence of two quasiholes. The effective anyon interactions defined by Eq.(10) with $N_{qh}=2$ are shown in Figure 3 (a) and (b), plotted against their relative angular momentum $\Delta M$. It depends on the parity of the system, as a manifestation of the odd-even effect in the MR phase [17]. Since the resulting dynamics have distinct behaviors under different PPs by which the e-e interaction can always be decomposed, it suggests that tailored electron-electron interactions could exploit this distinction to manipulate the two types of MR anyons. For instance, the e-e interaction dominated by $\hat{V}_{1}$ induces the effective interactions for $\psi$-type anyons to bind together while favoring a comparatively large structure of $1$-types. Conversely, to introduce an effectively repulsive interaction between the $\psi$-type anyons for non-abelian braiding (therefore the possibility of topological quantum computation), e-e interaction dominated by $\hat{V}_{3}$ is preferred.
@@
The resultant effective interactions between two localized $e/4$ MR quasiholes are plotted as a function of their chord distance. The odd-even effect remains prominent, with nearly opposite dynamical characteristics observed between systems with the two different parities. Under Coulomb interaction shown in Figure 3 (c), the effective interactions are repulsive at all distances for systems with $N_{e}$ of both parities. With shorter-range Yukawa interactions example $\lambda=0.5/l_{B}$ in Figure 3 (d)-the effective interaction for $\psi$-types is attractive at small separations. But it remains repulsive for $1$-types with the same distances, then develops to be attractive at a comparatively large separation. This behaviour indicates that short-range interactions energetically favor tightly bound pairs of $e/4$ MR anyons in odd $N_{e}$ systems, while promoting spatially extended, loose configurations of them in even $N_{e}$ systems. It means that the $\psi$ fusion channel is energetically more favorable under the short-range interaction than the $1$ channel when considering the fusion of two $\sigma$” anyons with $e/4$ charges. These results further support the feasibility of engineering bare electron-electron interactions to selectively separate and control the two types of fusion channels of the $e/4$ MR anyon and the consequent $e/2$ excitations.
@@ III.2 Interaction Within Four MR Anyon Clusters
We now look at a cluster of four MR quasiholes—each carrying a $e/4$ charge—which will be useful in Sec . IV. A four-MR-anyon cluster emerges upon the removal of a single electron from the ground state. The resulting system exhibits a degeneracy of $2^{4/2}-1=2$, where $1$- and $\psi$-types could coexist, adding significant complexity to the analysis. Specifically, the two pairs of quasiholes can only be two 1-types or two $\psi$-types in even $N_{e}$ systems, whereas the configuration necessarily consists of one 1-type and one $\psi$-type in odd $N_{e}$ systems [74]. The energy spectrum of this system under model electron-electron interactions $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ in the MR null space (in supplementary [70]) reveals two key features: a persistent odd-even effect in systems with four MR quasiholes and a preference of loose anyon configurations characterized by large relative angular momentum $\Delta M$ within the four-anyon cluster.
To extend our investigation to more realistic scenarios, given that MR states are proposed to be stabilized in the 1LL, we studied the e-e Coulomb and Yukawa potentials in the 1LL shown in Figure 4. In contrast to Laughlin’s case, the ground states of MR systems with four anyons consistently exhibit a loose structure and remain largely insensitive to variations in electron-electron interactions. For the system with $N_{e}=13$, the lowest energy state is the one with relative angular momentum $\Delta M=12$, the maximum allowed $\Delta M$ value in this finite system. In the $N_{e}=14$ case, the four-anyon cluster attains the largest possible value of $\Delta M=14$ under the 1LL electron-electron Coulomb interaction, and changes to $\Delta M=12$ for both Yukawa interactions we computed with $\lambda=0.5$ and $1$. The results are not conclusive due to the limited system size we can reach. Another observation is that the quasihole excitation spectrum maintains a comparatively robust, model-independent structure, unchanged not only across the parameter regimes studied among the realistic interactions, but also across the realistic interaction and the model Hamiltonian inside the MR CHS, whose spectrum can be further found in Supplementary Materials. Using the first and second moments of the density distribution, the estimated real-space radius of the four-quasihole cluster is $100\sim 140\text{nm}$ for the lowest energy state as $\Delta M=12$ in $N_{e}=15$ (odd) system. And for the $\Delta M=14$ anyon cluster in $N_{e}=16$ (even) system, it is $75\sim 110\text{nm}$. The spatial extent of the anyon clusters with the lowest energy in the MR phase exhibits a difference between odd and even electron systems, suggesting a parity effect in the cluster configuration in the finite systems. And the predicted spatial extents are much larger compared to Laughlin’s cases  [70].
@@ IV Bulk tunneling experiments
The effective interactions within anyon clusters have direct experimental relevance in recently developed STM measurements, whose detailed setups and implementations, both theory and experiment, have been discussed in several recent works [46, 49, 48]. In such experiments, single-electron tunneling occurs when the electronic potential $V_{t}$ applied to the tip satisfies
@@
where $E_{0}$ is the work function of the 2DEG, $E_{z}$ accounts for the finite-thickness effect [75], and $E_{xy}$ represents the energy scale of the strongly correlated 2D system studied here. After the bulk tunneling, a few-anyon cluster with total positive charge of +e+e will be left behind. The number of anyons in this cluster is determined by the specific FQH phase (e.g. $3$ for Laughlin 1/3 and $4$ for MR). Due to the non-trivial inter-anyonic interactions, as we have shown in the previous sections, the anyon cluster could be in different forms of “bound states” depending on the effective anyon interaction, the size of the STM tip, and the bias voltage applied to the tip. In particular, the effective interactions and the consequent structures of the few-anyon clusters influence the experimental outcomes, especially for the low-lying parts, which are mainly contributed by the quasihole states. It makes such measurements potentially very useful for understanding the dynamics of anyon clusters.
@@
Figure 5: Possible low-energy LDOS for Laughlin 1/3 state with $N_{e}=10$ at FWHM = $0.003\text{(first panel)}/0.005\text{(second panel)}\ e^{2}/\epsilon l_{B}$. The entire background energies are subtracted in the subfigures since only the energy differences matter here. (a) Coulomb interaction, (b) Yukawa interaction when $\lambda=1$, both in the LLL.
@@
The local density of states (LDOS) is a key observable in STM measurements. High-resolution tunneling can reveal a clear gap of approximately $0.1meV$, which corresponds to $\Delta E=0.002\sim 0.006\ e^{2}/\epsilon l_{B}$ whose exact value depends on the specific dielectric constant $\epsilon$ and the magnetic field $B$ under detection. According to this range, we simulated the possible LDOS for the 3-anyon Laughlin 1/3 states and 4-anyon MR states, focusing on the low-lying parts of the states mainly from the quasihole manifold. It is found that the variations in effective anyon interactions—induced by different bare e-e interactions—lead to distinct LDOS patterns, including differences in the numbers and positions of the peaks.
In this work, we follow the standard STM approximation where the tunneling amplitude is assumed to be constant and featureless, so the measured tunneling conductance is proportional to the local density of states (LDOS) of the sample. This approach omits the energy-dependent tunneling matrix element computed in more microscopic treatments such as Bardeen’s formalism, which derives from Fermi’s golden rule[76]. While the detailed tunneling amplitudes may vary with the STM tip geometry, material-specific parameters, or symmetry constraints, our analysis aims to capture the qualitative LDOS features arising from the internal structure of few-anyon bound states. A more refined modeling of the tunneling matrix elements will be pursued in future work.
@@
Our analysis shows that the low-lying tunneling spectrum is rich, affected by the effective anyon interaction and the internal structure of the anyon cluster, both in the Laughlin and Moore-Read phases. We only present a preliminary study here, assuming the bias voltage of the STM tip does not alter the anyon clusters’ densities. This may not be the case given that the one-body potential for an electron to tunnel highly depends on the geometry of the tip, the bias voltage, as well as the distance between the tip and the Hall manifold [74]. A more detailed study with explicit experimental parameters will be presented in the future.
@@ IV.1 LDOS Measurement for Laughlin 1/3 Phase
The low-energy spectrum for a system containing three Laughlin 1/3 anyons with realistic electron-electron interactions has already been shown in Figure 2 (a), where the most tightly bound anyon cluster state ($\Delta M=0$ state) lies at significantly higher energy than the other low-lying excitations for relatively long-range e-e interactions such as LLL Coulomb. As the screening parameter $\lambda$ increases, all quasihole excitations shift downward and gradually separate from the high-energy continuum. The $\Delta M=0$ state eventually becomes distinguishable from the gapped excitations but remains the highest in energy among all quasihole states. The simulated LDOS for $FWHM=0.003$ and $0.005$ under e-e Coulomb interaction and $\lambda=0.5$ Yukawa interaction, both in LLL, are shown in Figure 5. The LDOS results for different interactions are clearly distinguishable as in Figure 5. This indicates that LDOS can serve as a fingerprint for characterising the effective interactions among anyons.
Figure 6: LDOS of the MR system containing 4 anyons at FWHM = $0.001\ e^{2}/\epsilon l_{B}$ for: 1LL Coulomb interaction with (a1) $N_{e}=13$ and (a2) $N_{e}=14$; 1LL Yukawa interaction when $\lambda=0.5$ with (b1) $N_{e}=13$ and (b2) $N_{e}=14$.